# Time-dependent condensation of bosonic dysprosium

### v1

@@ 4 Ultracold dipolar dysprosium
Figure 2: Analytical vs. numerical NBDE solution for 164Dy. The absolute errors of the numerical NBDE solutions frelative to the analytical solutions shown in Fig. 1 are displayed for different times. They are defined as the absolute difference between the analytical and the numerical solution $|n_{\mathrm{ana}}(\epsilon,t)-n_{\mathrm{num}}(\epsilon,t)|.$
@@ 5 Time-dependent condensate fraction
Figure 4: Time-dependent condensate fraction for an equilibrating 3D 164Dy Bose gas. The calculations are based on the analytical solutions of the NBDE (see text) and displayed as functions of time for $a_{\mathrm{tot}}/a_{0}\equiv\sqrt{\sigma_{\mathrm{el}}}/a_{0}\simeq 637.90$, (solid curve, with $a_{s}/a_{0}=113$). Transport coefficients have then been recalculated for a reduced contact-interaction scattering length $a_{\mathrm{s}}/a_{0}=40$ (dot-dashed curve), and the correspondingly reduced total scattering length $a_{\mathrm{tot}}/a_{0}=355.30$ (dashed curve). The significant difference in scaling of the time-dependent condensate fraction with $a_{s}$ versus $a_{\mathrm{tot}}$ is expected to be measurable. It indicates the coherence mechanism (see text).
@@
The equilibrium value of the condensate fraction is determined by the initial particle number $N_{\mathrm{i}}(t)$ and the final temperature $T_{\mathrm{f}}$ as
@@
In this calculation, the solutions of the NBDE for constant coefficients would show an overshoot of the time-dependent condensate fraction above the equilibrium value $N^{\mathrm{eq}}_{\mathrm{c}}/N_{\mathrm{i}}$, before reaching that value from above. This is a consequence of the fact that the time-dependent solution of the NBDE for constant coefficients equilibrates too slowly in the UV. To avoid the overshoot, we calculate the distribution function from the exact NBDE solution ${n}(\epsilon,t)$ up to $\epsilon=\epsilon_{c}$ with $\epsilon_{c}=27.5$ nK for 39K and $t=\tau_{\mathrm{eq}}$, but use an approach to the equilibrium distribution according to a linear relaxation ansatz with time constant $\tau_{\mathrm{rel}}=90$ ms for $\epsilon>\epsilon_{c}$ and $t>\tau_{\mathrm{eq}}$. This is in line with the expectation that the bosonic enhancement – which is implicit in the nonlinear diffusion equation – becomes less relevant at higher energies, such that a relaxation ansatz for the time evolution with a short relaxation time $\tau_{\mathrm{rel}}$ can be taken. With this modification, the condensate fraction is found to be in agreement with the data, Fig. 3. In a parabolic trap with a steeper energy dependence of the density of states $\propto\epsilon^{2}$ as in our present 164Dy calculation, no such modification for constant transport coefficients is required to avoid an overshoot of the condensate fraction above the equilibrium value. The agreement with the 39K data confirms, in particular, the proportionality of the transport coefficients to the contact-interaction scattering lengths $a_{s}$, and the dependence of the time scales on $1/a_{s}$, rather than $1/a_{s}^{2}$. This indirect proof of emerging coherence in the formation of the condensate in alkali atoms has already been presented in Fig. 3 of Ref. [7].
@@
Our results should thus be taken as predictions, to be compared with forthcoming time-dependent condensate-formation data in dipolar atoms as functions of $a_{\mathrm{tot}}$ or $a_{\mathrm{s}}$ for different contact-interaction scattering lengths, as tuned by an external magnetic field. There is a clear and measurable difference between our predictions for a scaling with contact-interaction scattering lengths only, versus scaling with the effective total scattering lengths including the dipolar interaction, and experiments should decide between these two limiting possibilities. The difference determines whether coherence in time-dependent condensate formation is being built up only through the contact interaction, or mediated as well by the medium-range magnetic dipole interaction that includes higher partial waves.
It would also be of interest to investigate the 2D problem, with dipoles aligned in versus out-of-plane. In this case it is, however, not possible to average over the incident angles, as done here for the 3D situation, such that the present model must be extended accordingly.

### v2

@@ 4 Ultracold dipolar dysprosium
Figure 2: Analytical vs. numerical NBDE solution for 164Dy. The absolute errors of the numerical NBDE solutions relative to the analytical solutions shown in Fig. 1 are displayed for different times. They are defined as the absolute difference between the analytical and the numerical solution $|n_{\mathrm{ana}}(\epsilon,t)-n_{\mathrm{num}}(\epsilon,t)|.$
@@ 5 Time-dependent condensate fraction
Figure 4: Time-dependent condensate fraction for an equilibrating 3D 164Dy Bose gas. The calculations are based on the analytical solutions of the NBDE (see text) and displayed as functions of time for $a_{\mathrm{tot}}/a_{0}\equiv\sqrt{\sigma_{\mathrm{el}}}/a_{0}\simeq 637.90$, (solid curve, with $a_{s}/a_{0}=113$). Transport coefficients have then been recalculated for a reduced contact-interaction scattering length $a_{\mathrm{s}}/a_{0}=40$ (dot-dashed curve), and the correspondingly reduced total scattering length $a_{\mathrm{tot}}/a_{0}=355.30$ (dashed curve). The significant difference in scaling of the time-dependent condensate fraction with $a_{s}$ vs. $a_{\mathrm{tot}}$ is expected to be measurable. It indicates the coherence mechanism (see text).
@@
The equilibrium value of the condensate fraction is determined by the initial particle number $N_{\mathrm{i}}$ and the final temperature $T_{\mathrm{f}}$ as
@@
In this calculation, the solutions of the NBDE for constant coefficients would show an overshoot of the time-dependent condensate fraction above the equilibrium value $N^{\mathrm{eq}}_{\mathrm{c}}/N_{\mathrm{i}}$, before reaching that value from above. This is a consequence of the fact that the time-dependent solution of the NBDE for constant coefficients equilibrates too slowly in the UV. To avoid the overshoot, we calculate the distribution function from the exact NBDE solution ${n}(\epsilon,t)$ up to $\epsilon=\epsilon_{c}$ with $\epsilon_{c}=27.5$ nK for 39K and $t=\tau_{\mathrm{eq}}$, but use an approach to the equilibrium distribution according to a linear relaxation ansatz with time constant $\tau_{\mathrm{rel}}=90$ ms for $\epsilon>\epsilon_{c}$ and $t>\tau_{\mathrm{eq}}$. This is in line with the expectation that bosonic enhancement and condensate formation – which are implicit in the nonlinear diffusion equation – become less relevant at higher energies, such that a relaxation ansatz for the time evolution with a short relaxation time $\tau_{\mathrm{rel}}$ can be taken.
With this modification, the condensate fraction is found to be in agreement with the data, Fig. 3. In a parabolic trap with a steeper energy dependence of the density of states $\propto\epsilon^{2}$ as in our present 164Dy calculation, no such modification for constant transport coefficients is required to avoid an overshoot of the condensate fraction above the equilibrium value. The agreement with the 39K data confirms, in particular, the proportionality of the transport coefficients to the contact-interaction scattering lengths $a_{s}$, and the dependence of the time scales on $1/a_{s}$, rather than $1/a_{s}^{2}$. This indirect proof of emerging coherence in the formation of the condensate in alkali atoms has already been presented in Fig. 3 of Ref. [7].
@@
Our results should thus be taken as predictions, to be compared with forthcoming time-dependent condensate-formation data in dipolar atoms as functions of $a_{\mathrm{tot}}$ or $a_{\mathrm{s}}$ for different contact-interaction scattering lengths, as tuned by an external magnetic field. There is a clear and measurable difference between our predictions for a scaling with contact-interaction scattering lengths only, vs. scaling with the effective total scattering lengths including the dipolar interaction, and experiments should decide between these two limiting possibilities. The difference determines whether coherence in time-dependent condensate formation is being built up only through the contact interaction, or mediated as well by the medium-range magnetic dipole interaction that includes higher partial waves.
It would also be of interest to investigate the 2D problem, with dipoles aligned in vs. out-of-plane. In this case it is, however, not possible to average over the incident angles, as done here for the 3D situation, such that the present model must be extended accordingly.